# suppSupplementary References

### v1

@@ Abstract
Vision State Space Models (VSSMs), a novel architecture that combines the strengths of recurrent neural networks and latent variable models, have demonstrated remarkable performance in visual perception tasks by efficiently capturing long-range dependencies and modeling complex visual dynamics. However, their robustness under natural and adversarial perturbations remains a critical concern. In this work, we present a comprehensive evaluation of VSSMs’ robustness under various perturbation scenarios, including occlusions, image structure, common corruptions, and adversarial attacks, and compare their performance to well-established architectures such as transformers and Convolutional Neural Networks. Furthermore, we investigate the resilience of VSSMs to object-background compositional changes on sophisticated benchmarks designed to test model performance in complex visual scenes. We also assess their robustness on object detection and segmentation tasks using corrupted datasets that mimic real-world scenarios. To gain a deeper understanding of VSSMs’ adversarial robustness, we conduct a frequency analysis of adversarial attacks, evaluating their performance against low-frequency and high-frequency perturbations. Our findings highlight the strengths and limitations of VSSMs in handling complex visual corruptions, offering valuable insights for future research and improvements in this promising field. Our code and models will be available at https://github.com/HashmatShadab/MambaRobustness.
@@ 1 Introduction
VSSMs, such as the Vision Mamba [14] and VMamba [28] models, have gained attention in the vision domain due to their impressive performance. These models offer a unique approach to managing spatial dependencies, which is critical for handling dynamic visual environments. Their ability to selectively adjust interactions between states promises enhanced adaptability, a trait that could be pivotal in improving resilience against perturbations. Given their potential in safety-critical applications such as autonomous vehicles, robotics, and healthcare, it is crucial to thoroughly assess the robustness of these models.
@@
  • Occlusions and Information Loss: We rigorously assess the robustness of VSSMs against information loss along scanning directions and severe occlusions affecting foreground objects, non-salient background regions, and random patch drops at multiple levels. This analysis is crucial for understanding how well VSSMs can handle partial information drop and maintain performance despite occlusions. Additionally, we explore the sensitivity of VSSMs to the overall image structure and global composition through patch shuffling experiments, providing insights into their ability to capture and utilize global context.
@@
  • Common Corruptions: We evaluate the robustness of VSSM-based classification models against common corruptions that mimic real-world scenarios. This includes both global corruptions such as noise, blur, weather, and digital-based corruptions at multiple intensity levels, and fine-grained corruptions like object attribute editing and background manipulations. Furthermore, we extend the evaluation to VSSM-based detection and segmentation models to demonstrate their robustness in dense prediction tasks. By testing the models under these diverse and challenging conditions, we aim to provide a comprehensive understanding of their performance and resilience in real-world applications.
@@
Our findings highlight the strengths and limitations of VSSM models in handling various nuisances and adversarial attacks. While VSSMs demonstrate superior robustness in many scenarios, there are areas where other architectures, such as ConvNext and ViT, exhibit competitive or even better performance. These insights can guide researchers and practitioners in selecting the most appropriate model for their specific applications, taking into account the potential challenges and requirements for robustness in real-world settings.
@@ 2 Related Work
Robustness of Deep Learning Models: Robustness refers to a conventionally trained model’s ability to maintain satisfactory performance under natural and adversarial distribution shifts [12, 10]. In practice, deep learning-based models often encounter various types of corruptions, such as noise, blur, compression artifacts, and adversarial perturbations, which can significantly degrade their performance. To ensure the reliability and robustness of these models, it is essential to systematically evaluate their performance under such challenging conditions. Recent studies have investigated the robustness of deep learning-based models across a wide range of areas, including image classification [17, 19], semantic segmentation [21], object detection [36], video classification [45], point cloud processing [20], and transformer-based architectures [39, 24, 40]. However, there is a lack of similar investigations for vision state space models (VSSMs), despite their growing popularity and potential applications [27, 46, 44, 6]. In this work, we aim to bridge this gap by examining how the performance of VSSMs is affected by adversarial and common corruptions, thereby contributing to a more comprehensive understanding of model robustness across different architectural paradigms. Considering the increasing adoption of VSSMs, our findings can provide valuable insights for researchers and practitioners working on developing robust and reliable vision systems.
@@
3 Robustness Evaluation of Vision State Space Models
We have broadly categorized the experiments into natural and adversarial corruption categories to evaluate the robustness of CNNs, transformers, and VSSMs across tiny, small, and base-model families. For natural corruptions, we conduct experiments on classification, detection, and segmentation tasks. For the classification task, we employ the recent ConvNext model [30] as a representative of the CNN family, while selecting the ViT [9] and Swin [29] family models for transformer architectures. For VSSMs, we report results on VMamba v2 pretrained models [28]. For detection and segmentation, we report results using ImageNet-pretrained backbones of the specified models. These models are fine-tuned with the MMDetection [5] and MMSegmentation [7] frameworks. For detection, we utilize Mask-RCNN [16], and for segmentation, we use UperNet [43] as the network architecture. To evaluate the robustness of VSSMs against adversarial attacks, we consider imagenet trained classification models. Furthermore, we use imagenet pretrained models for adversarial fine-tuning on two downstream datasets; CIFAR-10 [22] and Imagenette [1]. All the evaluations for the task of classification are done on image size of $224\times 224$, except for CIFAR-10 where image size of $32\times 32$ is used. For detection and segmentation, evaluation is done on image size of $800\times 1216$ and $512\times 512$, respectively.
@@ 3.1 Robustness against Natural Corruptions
We categorize natural corruptions into information drop and ImageNet-based corruption benchmarks. Information drop experiments assess models’ robustness against various patch perturbations, such as occlusions, random patch drops, and patch shuffling, assessing their ability to handle partial information loss and local distortions. All information drop experiments are conducted on 5000 images from the ImageNet validation set, following [38]. ImageNet-based corruptions, on the other hand, are standard benchmarks that mimic real-world corruptions, including noise, blur, weather, background change, and digital-based corruptions at multiple intensity levels. We also demonstrate the robustness of VSSM-based detectors and segmentation models under natural corruptions.
Figure 1: Information drop of Tiny and Small family of models along the scanning direction: the image is split into a sequence of fixed-size non-overlapping patches of size 16x16, 8x8. 4x4, and 2x2. The top two rows show the results of linearly increasing the number of pixels dropped from each patch to the maximum threshold (Drop Intensity) along the scanning direction. The bottom two rows present results of linearly increasing the number of pixels dropped from each patch to the maximum threshold (Drop Intensity) until the center of the scanning direction. The plots on the left show results for top-to-bottom direction, while those on the right show results for left-to-right direction. More detailed analysis is provided in Section B.1 of Appendix.
@@ 3.1.1 Robustness against Information Drop
Information Drop along the Scanning Axis: VSSM models scan through input image patches sequentially along four distinct paths in order to capture the spatial information present in the image. These four paths comprise of scanning input patches in the directions of top to bottom, bottom to top, left to right, and right to left. To study the effectiveness of 2D-Selective Scan operation of VSSM model in aggregating the spatial information, we investigate the models’ behavior when information is dropped along the scanning directions. We consider two settings: (1) linearly increasing the amount of information dropped in each patch along the scanning direction, with the most information dropped in patches that are traversed last by the scanning operation, and (2) linearly increasing the amount of information dropped in patches from the start, reaching a maximum at the center, and then linearly decreasing the amount of information dropped until the end.
@@
Random Patch Drop: We assess the robustness of VSSMs in occluded scenarios by randomly dropping patches from images. We split the image into $n\times n$ patches and randomly select patches whose values will be set to zero. As shown in Tab.1 (top), when image is split into $16\times 16$ patches, VSSMs consistently outperform ResNet, ConvNeXt, and ViT models in maintaining accuracy with increasing numbers of dropped patches. However, under conditions of extreme spatial information loss, Swin models demonstrate superior performance. This trend persists when the image is split into $8\times 8$ patches, as illustrated in Tab.1 (bottom), highlighting the robustness of VSSMs and the exceptional resilience of Swin models in extreme patch drop scenarios. Results on different patch sizes showing similar trend are reported in Appendix B.4.
@@
VSSM-T
VSSM-S
VSSM-B
@@
VSSM-T
VSSM-S
VSSM-B
@@
Patch Shuffling: VSSMs process images as a sequence of patches, and the order of these patches represents the overall image structure and global composition. To evaluate the robustness of VSSMs to patch permutations, we define a shuffling operation on the input image patches, which destroys the image structure by changing the order of the patch sequence. Based on the dimensions of the patch size, we split the image into either $4,8,16,32,64$ and $256$ patches, which is then followed by random shuffling of the patches to evaluate the performance of models on spatial rearrangement of information. In Tab.  2 (bottom), we observe that VSSM family overall performs better than other models when the spatial structure of the input image is disturbed. VSSM models generally demonstrate greater resilience to spatial structure disturbances than Swin models.
Corruptions $\downarrow$
ResNet-50
ConvNext-T
ConvNext-S
ConvNext-B
ViT-T
ViT-S
ViT-B
VSSM-T
VSSM-S
VSSM-B
Swin-T
Swin-S
Swin-B
Brightness 100.00 68.20 58.63 52.15 223.14 101.09 41.67 61.60 53.06 52.94 75.89 65.35 63.54
Contrast 100.00 40.14 30.29 28.96 426.29 186.49 103.16 38.17 30.02 27.70 44.41 42.09 41.70
Defocus Blur 100.00 85.70 77.82 73.87 103.27 70.64 42.81 84.96 72.32 73.00 87.50 66.82 81.49
Elastic Transform 100.00 81.11 72.18 69.44 70.34 55.69 41.43 82.78 69.89 72.21 86.36 66.53 76.78
Fog 100.00 113.68 86.18 91.19 312.26 119.54 68.79 87.47 64.93 55.88 84.32 87.24 68.61
Frost 100.00 64.12 57.09 50.75 146.53 75.52 27.19 42.05 36.58 30.79 47.77 40.72 38.63
Gaussian Blur 100.00 86.69 84.58 77.75 107.84 82.50 55.22 88.91 77.05 77.21 90.59 81.36 85.37
Gaussian Noise 100.00 52.01 51.36 56.02 130.74 66.75 30.16 56.81 38.29 52.62 73.31 59.86 54.49
Glass Blur 100.00 89.71 85.46 81.43 98.30 84.73 66.52 93.59 86.00 87.28 96.53 86.27 90.24
Impulse Noise 100.00 47.18 44.71 45.78 122.95 61.19 28.27 44.55 31.93 39.72 63.83 48.63 48.68
JPEG Compression 100.00 69.79 59.62 60.96 131.33 78.90 43.49 82.29 65.82 63.30 144.27 88.13 108.22
Motion Blur 100.00 65.05 56.10 51.42 106.50 58.86 33.26 65.84 54.55 54.13 80.88 50.12 62.42
Pixelate 100.00 82.35 71.52 66.00 52.11 27.31 15.88 80.34 66.30 65.02 109.24 100.70 101.51
Saturate 100.00 63.44 53.74 47.61 250.68 116.28 48.06 57.26 47.97 47.07 75.29 62.40 58.34
Shot Noise 100.00 52.69 51.85 56.73 128.00 66.02 30.12 56.20 38.75 51.72 73.51 62.09 57.99
Snow 100.00 54.55 48.57 43.18 134.88 61.12 26.04 54.29 46.02 42.14 57.70 39.38 47.00
Spatter 100.00 39.86 32.09 29.41 117.51 48.16 21.03 33.10 28.59 24.33 32.71 23.79 22.92
Speckle Noise 100.00 44.78 40.91 34.08 132.49 60.64 26.89 42.92 30.43 34.07 61.92 48.85 48.80
Zoom Blur 100.00 82.47 68.22 64.74 125.18 84.89 52.90 82.72 70.04 70.37 93.27 77.34 77.96
mCE 100.00 67.55 59.52 56.92 153.7 79.28 42.26 65.05 53.08 53.76 77.86 63.04 64.98
@@ 3.1.2 Robustness against ImageNet Corruptions
To evaluate the robustness of VSSMs in real-world scenarios, we performed experiments on several corrupted versions of the ImageNet dataset, which can be grouped into two categories based on the type of changes they introduce to the images. The first category includes datasets that make overall global compositional changes, such as ImageNet-C [17], which evaluates robustness against 19 common distortions across five categories (noise, blur, weather, and digital-based corruptions) with varying severity levels. Additionally, datasets like ImageNetV2 [41], ImageNet-A [19], ImageNet-R [18], and ImageNet-S [11] also fall into this category, as they introduce domain shifts that affect the global composition of the images. For details of these datasets, see Appendix B.5. The second category consists of datasets that provide more control in editing the images and focus on fine-grained details. ImageNet-E [26] assesses classifiers’ robustness to changes in background, object size, position, and direction, while ImageNet-B [34] introduces diverse object-to-background changes using text-to-image, image-to-text, and image-to-segment models, preserving original object semantics while varying backgrounds to include both natural and adversarial changes. These datasets use diffusion models to generate various object-to-background compositional changes to test the resilience of models.
@@
VSSM-T 82.28 71.16(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-11.12)}}}$ 33.99(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-48.29)}}}$ 12.08(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-70.20)}}}$ 32.05(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-50.23)}}}$ 37.32(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.96)}}}$
VSSM-S 83.48 73.01(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.47)}}}$ 36.98(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-46.50)}}}$ 16.45(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-67.03)}}}$ 35.75(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-47.73)}}}$ 40.55(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.93)}}}$
VSSM-B 83.76 73.22(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-10.54)}}}$ 38.53(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-45.23)}}}$ 18.35(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-65.41)}}}$ 35.99(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-47.77)}}}$ 41.52(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.24)}}}$
@@
VSSM-T 91.15 89.87 75.18 87.41 92.09 91.06 83.66 79.71 95.84 98.50 92.20 96.30 87.20 86.80
VSSM-S 92.03 90.79 76.15 88.81 93.22 92.25 85.57 81.89 96.37 99.20 94.10 97.40 90.90 89.50
VSSM-B 92.37 91.27 77.30 89.11 93.70 92.64 86.03 83.62 96.37 99.10 94.00 96.50 90.80 89.80
@@
Robustness against Global Corruptions: The performance of various models on ImageNet-C is reported in terms of the mean corruption error (mCE) in Tab. 3. The mCE (lower is better) represents the average error of the model across various common corruptions at multiple intensity levels, normalized by the ResNet-50’s standard accuracy. VSSM-T has the lowest mCE among all the ‘T’ versions, followed by VSSM-S, which performs significantly better than its ‘S’ counterparts. However, among the ‘B’ family, ViT-B achieves the lowest mCE score, possibly due to its pretraining on ImageNet21k and finetuning on ImageNet 1k. Notably, the mCE error of all Base models (VSSM, Swin, ConvNext) is larger than their Small counterparts, indicating that robust performance may not uniformly improve with model size. Similarly, the top-1 classification accuracy on domain-shifted datasets, including ImageNetV2 [41], ImageNet-A [19], ImageNet-R [18], and ImageNet-S [11] is shown in Tab. 4. VSSM-T and VSSM-S have the least average performance drop compared to their Swin and ConvNext-based counterparts. For Base models, ViT performs the best, which may be attributed to its pretraining on the larger ImageNet21k dataset.
@@
Robustness against Fine-grained Corruptions: Tab. 5 (left) shows the robustness of models on ImageNet-E against variations in background complexity, object size, and object positioning. Lower $\lambda$ values correspond to low background texture complexity, whereas higher $\lambda$ values indicate high background texture complexity, with $\lambda=20$ (adv) representing the extreme case of adversarially optimized high texture complexity. The object in the scene is resized proportionally to its original size, specifically to $(0.1,0.08,0.05)$, and is placed at random locations within the scene. As can be observed from the results, VSSM models show higher resilience to background changes compared to the corresponding Swin and ViT transformer families, with performance comparable to the advanced ConvNeXt models. As background complexity increases or object size decreases, the performance of all models declines, but the drop is less significant for the VSSM and ConvNeXt models compared to the Transformer-based models, demonstrating their robustness to object size variations. Similar trends are observed for the ImageNet-B dataset in Tab. 5 (right). This dataset introduces background changes in the images by using a diffusion model to modify the background with textual guidance, such as class or caption information as prompts, or by adding color and texture prompts. The VSSM family demonstrates superior performance compared to all Transformer-based variants and maintains performance better or comparable to the advanced ConvNext models. For further analysis across other model families, refer to Appendix B.5.
3.1.3 Robustness Evaluation for Object Detection

### v2

@@ Abstract
Vision State Space Models (VSSMs), a novel architecture that combines the strengths of recurrent neural networks and latent variable models, have demonstrated remarkable performance in visual perception tasks by efficiently capturing long-range dependencies and modeling complex visual dynamics. However, their robustness under natural and adversarial perturbations remains a critical concern. In this work, we present a comprehensive evaluation of VSSMs’ robustness under various perturbation scenarios, including occlusions, image structure, common corruptions, and adversarial attacks, and compare their performance to well-established architectures such as transformers and Convolutional Neural Networks. Furthermore, we investigate the resilience of VSSMs to object-background compositional changes on sophisticated benchmarks designed to test model performance in complex visual scenes. We also assess their robustness on object detection and segmentation tasks using corrupted datasets that mimic real-world scenarios. To gain a deeper understanding of VSSMs’ adversarial robustness, we conduct a frequency-based analysis of adversarial attacks, evaluating their performance against low-frequency and high-frequency perturbations. Our findings highlight the strengths and limitations of VSSMs in handling complex visual corruptions, offering valuable insights for future research. Our code and models will be available at https://github.com/HashmatShadab/MambaRobustness.
@@ 1 Introduction
VSSMs, such as the VMamba [29] and the hybrid Mamba-Transformer variant MambaVision [16], have gained attention in the vision domain due to their impressive performance. These models offer a unique approach to managing spatial dependencies, which is critical for handling dynamic visual environments. Their ability to selectively adjust interactions between states promises enhanced adaptability, a trait that could be pivotal in improving resilience against perturbations. Given their potential in safety-critical applications such as autonomous vehicles, robotics, and healthcare, it is crucial to thoroughly assess the robustness of these models.
@@
  • Occlusions and Information Loss: We rigorously assess the robustness of pure and hybrid VSSMs against information loss along scanning directions and severe occlusions affecting foreground objects, non-salient background regions, and random patch drops at multiple levels. This analysis is crucial for understanding how well VSSMs can handle partial information drop and maintain performance despite occlusions. Additionally, we explore the sensitivity of VSSMs to the overall image structure and global composition through patch shuffling experiments, providing insights into their ability to capture global context.
@@
  • Common Corruptions: We evaluate the robustness of VSSM-based classification models against common corruptions that mimic real-world scenarios. This includes both global corruptions such as noise, blur, weather, and digital-based corruptions at multiple intensity levels, and fine-grained corruptions like object attribute editing and background manipulations. Furthermore, we extend the evaluation to VSSM-based detection and segmentation models to demonstrate their robustness in dense prediction tasks. By testing the models under these diverse and challenging conditions, we aim to provide a comprehensive understanding of their resilience in real-world applications.
@@
Our findings reveal that VSSM models have both strengths and limitations in handling various nuisances and adversarial attacks. They also indicate that a variety of metrics is essential to fully evaluate the diverse capabilities of different architectures. While VSSMs often demonstrate superior robustness, ConvNext and ViT architectures occasionally outperform them. These insights can inform model selection for specific applications, considering robustness requirements in real-world scenarios.
@@ 2 Related Work
Robustness of Deep Learning Models: Robustness refers to a conventionally trained model’s ability to maintain satisfactory performance under natural and adversarial distribution shifts [12, 10]. In practice, deep learning-based models often encounter various types of corruptions, such as noise, blur, compression artifacts, and adversarial perturbations, which can significantly degrade their performance. To ensure the reliability and robustness of these models, it is essential to systematically evaluate their performance under such challenging conditions. Recent studies have investigated the robustness of deep learning-based models across a wide range of areas, including image classification [19, 20], semantic segmentation [22], object detection [37], video classification [46], point cloud processing [21], and transformer-based architectures [40, 25, 41]. However, there is a lack of similar investigations for vision state space models (VSSMs), despite their growing popularity and potential applications [28, 47, 45, 5]. In this work, we aim to bridge this gap by examining how the performance of VSSMs is affected by adversarial and common corruptions.Considering the increasing adoption of VSSMs, our findings can provide valuable insights for researchers and practitioners working on developing robust and reliable vision systems.
@@
3 Robustness of Vision State Space Models
We have broadly categorized the experiments into natural and adversarial corruption categories to evaluate the robustness of CNNs, transformers, and VSSMs across tiny, small, and base-model families. For natural corruptions, we conduct experiments on classification, detection, and segmentation tasks. For the classification task, we employ the recent ConvNext model [31] as a representative of the CNN family, while selecting the ViT [9] and Swin [30] family models for transformer architectures. For VSSMs, we report results on pure VMamba v2 pretrained models [29] and hybrid MambaVision models [16]. For detection and segmentation, we report results using ImageNet-pretrained backbones of the specified models. These models are fine-tuned with the MMDetection [6] and MMSegmentation [7] frameworks. For detection, we utilize Mask-RCNN [17], and for segmentation, we use UperNet [44] as the network architecture. To evaluate the robustness of VSSMs against adversarial attacks, we consider imagenet trained classification models. Furthermore, we use imagenet pretrained models for adversarial fine-tuning on two downstream datasets; CIFAR-10 [23] and Imagenette [1]. Evaluations are done on 224×224 images for classification (except 32×32 for CIFAR-10), 800×1216 for detection, and 512×512 for segmentation.
@@ 3.1 Robustness against Natural Corruptions
We categorize natural corruptions into information drop and ImageNet-based corruption benchmarks. Information drop experiments assess models’ robustness against various patch perturbations, such as occlusions, random patch drops, and patch shuffling, assessing their ability to handle partial information loss and local distortions. All information drop experiments are conducted on 5000 images from the ImageNet validation set, following [38]. ImageNet-based corruptions mimic real-world issues such as noise, blur, weather, and digital corruptions at various intensities. We also evaluate VSSM-based detectors and segmentation models under these natural corruptions.
Figure 1: Information drop of Tiny and Small family of models along the scanning direction: the image is split into a sequence of fixed-size non-overlapping patches of size 16x16, 8x8, and 4x4. The first row shows the results of linearly increasing the number of pixels dropped from each patch to the maximum threshold (Drop Intensity) along different scanning directions. The bottom row presents results of linearly increasing the number of pixels dropped from each patch to the maximum threshold (Drop Intensity) until the center of the scanning direction. More detailed analysis is provided in Section A.1 of the Appendix.
@@ 3.1.1 Robustness against Information Drop
Information Drop along the Scanning Axis: VSSM models scan image patches sequentially along four paths (top-bottom, bottom-top, left-right, right-left) to capture spatial information. To study the effectiveness of this 2D-Selective Scan operation, we investigate the models’ response to information drop along these scanning directions. We consider two settings: (1) linearly increasing information drop along the scanning direction, with maximum drop in last-scanned patches, and (2) linearly increasing drop from start to the center, then linearly decreasing until the end.
@@
Random Patch Drop: We assess the robustness of VSSMs in occluded scenarios by randomly dropping patches from images. We split the image into $n\times n$ patches and randomly select the patches whose values will be set to zero. As shown in Tab.1 (top), when image is split into $16\times 16$ patches, VSSMs consistently outperform MambaVision, ResNet, ConvNeXt, and ViT models in maintaining accuracy with increasing numbers of dropped patches. However, under conditions of extreme spatial information loss, Swin models demonstrate superior performance, whereas the hybrid architecture-based MambaVision performs worst. This trend persists when the image is split into $8\times 8$ patches, as illustrated in Tab.1 (bottom), highlighting the robustness of VSSMs and the exceptional resilience of Swin models in extreme patch drop scenarios. Results on different patch sizes showing similar trend are reported in Appendix A.2.
@@
VMamba-T
VMamba-S
VMamba-B
MambaVision-T
MambaVision-S
MambaVision-B
@@
VMamba-T
VMamba-S
VMamba-B
MambaVision-T
MambaVision-S
MambaVision-B
@@
Patch Shuffling: VSSMs process images as a sequence of patches, and the order of these patches represents the overall image structure and global composition. To evaluate the robustness of VSSMs to patch permutations, we define a shuffling operation on the input image patches, which destroys the image structure by changing the order of the patch sequence. Based on the dimensions of the patch size, we split the image into either $4,8,16,32,64$ and $256$ patches, which is then followed by random shuffling of the patches to evaluate the performance of models on spatial rearrangement of information. In Tab.  2 (bottom), we observe that VMamba family overall performs better than other models when the spatial structure of the input image is disturbed. VMamba models generally demonstrate greater resilience to spatial structure disturbances than Swin models.
@@ 3.1.2 Robustness against ImageNet Corruptions
To evaluate the robustness of VSSMs in real-world scenarios, we performed experiments on several corrupted versions of the ImageNet dataset, which can be grouped into two categories based on the type of changes they introduce to the images. The first category includes datasets that make overall global compositional changes, such as ImageNet-C [19], which evaluates robustness against 19 common distortions across five categories (noise, blur, weather, and digital-based corruptions) with varying severity levels. Additionally, datasets like ImageNetV2 [42], ImageNet-A [20], ImageNet-R [18], and ImageNet-S [11] also fall into this category, as they introduce domain shifts that affect the global composition of the images. For details of these datasets, see Appendix B.1. The second category consists of datasets that provide more control in editing the images and focus on fine-grained details. ImageNet-E [27] assesses classifiers’ robustness to changes in background, object size, position, and direction, while ImageNet-B [35] introduces diverse object-to-background changes using text-to-image, image-to-text, and image-to-segment models, preserving original object semantics while varying backgrounds to include both natural and adversarial changes. These datasets use diffusion models to generate various object-to-background compositional changes to test the resilience of the models.
@@
Model $\downarrow$ ImageNet OOD Average($\uparrow$) mCE($\downarrow$)
ConvNext-T 81.87 67.55 36.90(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.97)}}}$
ViT-T 75.35 153.7 26.88(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-48.47)}}}$
Swin-T 80.91 77.86 34.00(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-46.91)}}}$
VMamba-T 82.28 65.05 37.32(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.96)}}}$
MambaVision-T 82.10 66.58 37.66(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.44)}}}$
ConvNext-S 82.82 59.52 39.81(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-43.01)}}}$
ViT-S 81.40 79.28 v 36.74(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.66)}}}$
Swin-S 82.90 63.04 37.81(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-45.09)}}}$
VMamba-S 83.48 53.08 40.55(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.93)}}}$
MambaVision-S 83.22 50.50 39.52(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-43.70)}}}$
ConvNext-B 83.75 56.92 41.68(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.07)}}}$
ViT-B 84.40 42.26 45.49(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-38.91)}}}$
Swin-B 83.08 64.98 38.97(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-44.11)}}}$
VMamba-B 83.76 53.76 41.52(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-42.24)}}}$
MambaVision-B 83.96 49.61 41.98(-${}_{\text{{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}(-41.98)}}}$
@@
VMamba-T 91.15 89.87 75.18 87.41 92.09 91.06 83.66 79.71 95.84 98.50 92.20 96.30 87.20 86.80
VMamba-S 92.03 90.79 76.15 88.81 93.22 92.25 85.57 81.89 96.37 99.20 94.10 97.40 90.90 89.50
VMamba-B 92.37 91.27 77.30 89.11 93.70 92.64 86.03 83.62 96.37 99.10 94.00 96.50 90.80 89.80
MambaVision-T 90.67 88.83 73.07 86.40 91.93 90.07 81.87 78.42 95.73 98.60 93.70 96.60 89.10 87.70
MambaVision-S 91.22 90.19 75.18 88.19 92.78 91.19 84.01 81.18 96.03 99.40 94.40 97.80 91.40 90.10
MambaVision-B 91.77 90.65 78.42 89.18 93.70 92.81 86.35 83.78 96.30 99.10 94.60 97.20 91.40 90.60
@@
Robustness against Global Corruptions: The performance of various models on ImageNet-C is reported in terms of the mean corruption error (mCE) in Tab. 3. The mCE (lower is better) represents the average error of the model across various common corruptions at multiple intensity levels, normalized by the ResNet-50’s standard accuracy. VMamba-T has the lowest mCE among all the ‘T’ versions, followed by MambaVision-S, which performs significantly better than its ‘S’ counterparts. However, among the ‘B’ family, ViT-B achieves the lowest mCE score, , possibly due to ImageNet21k pretraining and ImageNet 1k finetuning. The table also shows top-1 accuracy on domain-shifted datasets (ImageNetV2, ImageNet-A, ImageNet-R, ImageNet-S). ‘T’ and ‘S’ versions of VMamba and MambaVision demonstrate the least average performance drop compared to Swin and ConvNext counterparts. For ‘B’ models, ViT performs best, likely due to ImageNet21k pretraining. For more results, see Appendix B.1. Furthermore, in Appendix B.2, we report Expected Calibration Error(ECE) of models across the mentioned datasets to quantify the reliability of model’s predicted confidence levels, accompanied by reliability diagrams for visualization.
@@
Robustness against Fine-grained Corruptions: Tab. 4 (left) shows model robustness on ImageNet-E against variations in background complexity, object size, and positioning. Lower $\lambda$ values indicate low background texture complexity, higher values indicate high complexity, with $\lambda=20$ (adv) representing adversarially optimized high texture complexity. Objects are resized to (0.1, 0.08, 0.05) of original size and randomly placed. VSSM models (VMamba and MambaVision) demonstrate higher resilience to background changes compared to Swin and ViT transformer families, and perform comparably to ConvNeXt models. As background complexity increases or object size decreases, all models’ performance declines, but the drop is less significant for VSSM and ConvNeXt models, showing their robustness to object size variations. Similar trends are observed for ImageNet-B (Tab. 4, right), where backgrounds are modified using a diffusion model with textual guidance (class/caption information) or color/texture prompts. The VSSM family, including VMamba and MambaVision, demonstrates superior performance compared to all Transformer-based variants and maintains performance better or comparable to the advanced ConvNext models. Pure VMamba and hybrid MambaVision exhibit comparable robustness against ImageNet corruption when no information is dropped. For further analysis across more models, see Appendix B.3.
3.1.3 Robustness on Object Detection